OpenStudy (donka74):

giving a gold medal solve system by substitution(explain) -5x-y=-3 3x-8y=24 how many solution? what is the nature system?

OpenStudy (anonymous):

take the first equation -5x-y=-3 y=-5x+3 now take the second equation 3x-8y=24 but y=-5x+3 so : 3x-8(-5x+3)=24 solve for x and use y=-5x+3 to find y